David drove home from college traveling an average speed of 55.4 mph and drove back to the college the following week at an average speed of 60.1 mph. If the total round trip took 11 hours, how much time did it take David to drive from home back to college? Express the time in hours and minutes. Round to the nearest minute.

Answer by stanbon(75887)

David drove home from college traveling an average speed of 55.4 mph and drove back to the college the following week at an average speed of 60.1 mph. If the total round trip took 11 hours, how much time did it take David to drive from home back to college? Express the time in hours and minutes. Round to the nearest minute.
--------
Going home DATA:
rate = 55.4 mph ; dist = x miles ; time = d/r = x/55.4 hrs
-------
Going to college DATA:
rate = 60.1 mph ; dist = x miles ; time = x/60.1 hrs
-----
Equation:
x/55.4 + x/60.1 = 11
-----
60.1x + 55.4x = 11*55.4*60.1
------
115.5x = 36624.94
x = 317.1 miles
------
Ans: home to college time = 317.1/60.1 = 5.28 hrs = 5 hrs 17 min
----------
Cheers,
Stan H.
----------